OpenStudy (anonymous):

Write an equation of the line that passes through (2,-1) and is perpendicular to the line 3y - x = 12 Medal will be given :)

OpenStudy (mertsj):

What is the slope of the given line?

OpenStudy (anonymous):

I dont know. I wrote down everything they gave me...

OpenStudy (mertsj):

Put the equation in the form y = mx+b

OpenStudy (anonymous):

y = x + 4

OpenStudy (mertsj):

You must divide EVERY term by 3

OpenStudy (anonymous):

y = x/3 + 4?

OpenStudy (mertsj):

So what is the slope...it is the coefficient of x

OpenStudy (mertsj):

I take it you're not really in to this problem.

OpenStudy (anonymous):

The slope is 1? And no, i am into this problem.

OpenStudy (mertsj):